rootbeer003:

Who can help me with this word problem. representing and solving it with using a graph, table and algebra?

rootbeer003:

You want to join a health club for one year. Club A charges $55 per month and has no initiation fee. Club B costs $45 a month and has an initiation fee of $100. At what month are the costs of the two clubs the same?

ThisGirlPretty:

@563blackghost

Ultrilliam:

@dude

dude:

Lets make an equation for Club A first A line is written as \(y=mx+b\) m is the rate, and b is the initial point Based on this can you make the equation for Club A?

rootbeer003:

y = 55x umm

dude:

Right

rootbeer003:

oh alright

dude:

`Club B costs $45 a month and has an initiation fee of $100` Now we do the same thing for Club B ____________________________________________________ \(y=mx+b\) A line is written as y=mx+b m is the rate, and b is the initial point ____________________________________________________

rootbeer003:

would u say the intuition fee is the initial point?

dude:

Realized I copy pasted that incorrectly but it will have to do.

dude:

Yes

rootbeer003:

y = 45x + 100

dude:

Looks good

dude:

Now is asks when they will be the same Another word for same is equal aka = Set them equal to each other

rootbeer003:

55x = 45x + 100

dude:

Right, not get x on one side of the equation

rootbeer003:

55x = 45x + 100 /45 /45 /45

rootbeer003:

1.2x = 2.2 ?

dude:

Not quite, opposite of +45x is subtracting 45x on both sides

rootbeer003:

OH

rootbeer003:

55x = 45x + 100 -45 -45 -45 10X = 55

dude:

You're subtracting 45x on both sides, 100 and 45x are not like terms so you cannot subtract them \(~~~~55x = 45x + 100\\ -45x ~-45x \\ \_\_\_\_\_\_\_\_\_\_\_\_\_\_\_\\ ~~~10x = 100\)

rootbeer003:

Oh, so only from the X's not all of them

rootbeer003:

someone told me all of them woopsie daisy

rootbeer003:

someone told me all of them woopsie daisy

rootbeer003:

10x =100 /10 /10

dude:

Yes

rootbeer003:

x = 10

dude:

Right x = months So after 10 months, they will have the same cost

rootbeer003:

oooooo

rootbeer003:

oooooo

rootbeer003:

so now i have to do this on a graph

dude:

You could use technology for this but are you graphing this manually?

rootbeer003:

i could use desmos, do i just apply the two formulas?

dude:

Yeah \(y=55x\) and \( y= 45x + 100\)

rootbeer003:

alright ill do that rn, in then lastly its doing a table

dude:

You can use the graph to find the points, just write x=1, x=2 and x=3 Where they intersect is the point

rootbeer003:

wait i did it wrong i think where did u get those numbers

rootbeer003:

1 attachment
dude:

That is right, the window is just out of place because you're only looking at single digits xD

rootbeer003:

oH but how do i find the points if they look like this

rootbeer003:

1 attachment
rootbeer003:

1 attachment
dude:

Yeh

rootbeer003:

what do i do wit the 550

dude:

That is just the cost, when they are both at 10 months

rootbeer003:

ooHHHHHHHHHHHHHHhhhh

rootbeer003:

ok now for the table :D:D:D:D:D:D

rootbeer003:

lmao

rootbeer003:

ill make the table n see if it looks right

dude:

Okay

rootbeer003:

1 attachment
rootbeer003:

hows this

rootbeer003:

should prob make the left part longer cuz its months?

rootbeer003:

1 attachment
dude:

Dude, every x value is a month, you only need 2 rows xD

rootbeer003:

wdf

rootbeer003:

ok ignore the left part

rootbeer003:

1 attachment
rootbeer003:

:3

dude:

It was right, rows are the horizontal xD You dont need many of the horizontal unless it was given to you like that

rootbeer003:

i thought i would make 10 slots for 10 months yknow

dude:

You can do that

dude:

Unless it specified, its up to you

rootbeer003:

ill do it if it works

rootbeer003:

so for the cost of club a, it doesnt have a rising point

dude:

Rising point?

rootbeer003:

OH I MEANT initial point

dude:

Yes, that starts at 0

rootbeer003:

so they both start at 0 on day 0?

dude:

No, just Club A starts at 0

rootbeer003:

1 attachment
dude:

Yes, the months for that would be 0

rootbeer003:

so the months start at 0 too?

dude:

Yeah because at 0 months, the cost is 0 for Club A

rootbeer003:

and club B starts with 100?

dude:

Yes

rootbeer003:

1 attachment
dude:

Right

rootbeer003:

ill do the club b rq

rootbeer003:

1 attachment
dude:

Looks good :P

rootbeer003:

so do i start club a with 55?

dude:

Month 1 is 55, yes

rootbeer003:

and i add 55 each time?

dude:

Yes

rootbeer003:

okie

rootbeer003:

1 attachment
rootbeer003:

BAM

dude:

Perfect

rootbeer003:

yayy thankssss

dude:

Ofcs